solve by using elimiation/addtion
2m - 5n = 13
5m + 3n = 17
show steps

Answer by mananth(16946)

2 m -5 n = 13 .............1
5 m + 3 n = 17 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 5
6 m -15 n = 39
25 m 15 n = 85
Add the two equations
31 m = 124
/ 31
m = 4.00
plug value of m in (1)
2 m + -5 n = 13
8 + -5 n = 13
-5 n = 13 -8
-5 n = 5
n = -1.00